# Fractional diffusion as the limit of a short range potential Rayleigh gas

### v1

@@ 3.2. Formula of the collision kernel $B$
where the first term is equal to a finite fixed number $a$, since $n$ is fixed, and the second term is equal to $\mathcal{O}(\frac{\lambda}{b})$. Therefore we have that
$\int_{0}^{1-\frac{\lambda}{b}}\frac{\mathrm{d}w}{\sqrt{1-(1-w)^{n-1}}}=a+\mathcal{O}(\frac{\lambda}{b}).$
@@
$\Sigma^{\prime}(0)=1$ $+\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}\int_{0}^{1-\frac{\lambda}{b}}\frac{\mathrm{d}w}{\sqrt{1-(1-w)^{n-1}}}=1+\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}\bigg{(}a+\mathcal{O}(\frac{\lambda}{b})\bigg{)}.$
@@
$\Sigma^{\prime}(0)=1+a\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}+\mathcal{O}(\frac{\lambda}{b})\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}.$ (3.12)
[…]
@@
$\int_{c}^{1}\frac{\mathrm{d}y}{\sqrt{1-y^{n-1}}}=\int_{0}^{1-c}\frac{\mathrm{d}z}{\sqrt{1-(1-z)^{n-1}}}=\int_{0}^{1-c}\frac{\mathrm{d}z}{\sqrt{(n-1)z}}=\frac{2(1-c)^{\frac{1}{2}}}{\sqrt{n-1}}<\infty,$
where in the first equality we used the change of variable $z=1-y$ and in the second equality we used Taylor expansion of first order of the function $(1-z)^{n-1}$ around $z_{0}=0$. Here, $c$ is a positive fixed number, less than $1$. The function $j_{\delta}(y)$ can be written as
@@
$\int_{\frac{\lambda}{b(1-\delta)}}^{1}\frac{\mathrm{d}y}{\sqrt{j_{\delta}(y)}}\to\int_{c}^{1}\frac{\mathrm{d}y}{\sqrt{1-y^{n-1}}}\quad\text{as}\ \delta\to 0.$
@@
$\frac{\lambda}{b}=\frac{\bigg{(}1+\frac{2K}{\mathcal{E}_{0}}\bigg{)}^{\frac{1}{2}}}{\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}}\approx\frac{1}{\mathcal{E}_{0}^{\frac{1}{n-1}}}.$
@@
$\frac{\lambda}{b}\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}\approx 1,\quad\textrm{i.e.}\quad\mathcal{O}(\frac{\lambda}{b})\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}=\mathcal{O}(1).$
Therefore
$\Sigma^{\prime}(0)=1+\frac{a}{(2K)^{\frac{1}{n-1}}}\bigg{(}2K^{2}+\mathcal{E}_{0}\bigg{)}^{\frac{1}{n-1}}+\mathcal{O}(1)$ $=1+a\bigg{[}\bigg{(}\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}+\frac{K}{n-1}\bigg{(}\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{2-n}{n-1}}\bigg{]}+\mathcal{O}(1)$
$=a\bigg{(}\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{1}{n-1}}+\underbrace{\frac{aK}{n-1}\bigg{(}\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{2-n}{n-1}}+\mathcal{O}(1)}_{=\mathcal{O}(1)},$
where in the second equality we used the Taylor expansion of first order of the function $h(x):=\big{(}x+\mathcal{E}_{0}\big{)}^{\frac{1}{n-1}}$, with $x:=2K^{2}$, around $x_{0}=0$. Thus, by the inverse function theorem
$(\Sigma^{-1})^{\prime}(\Sigma(0))=\frac{1}{a\big{(}\frac{\mathcal{E}_{0}}{2K}\big{)}^{\frac{1}{n-1}}+\mathcal{O}(1)}$ $=a^{-1}\bigg{(}\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{-\frac{1}{n-1}}\frac{1}{1+a^{-1}\big{(}\frac{\mathcal{E}_{0}}{2K}\big{)}^{-\frac{1}{n-1}}\mathcal{O}(1)}$
$=a^{-1}\bigg{(}\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{-\frac{1}{n-1}}\big{(}1+\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}})\big{)},$
@@
$(\Sigma^{-1})^{\prime}(\Sigma(0))=\frac{1}{a}(2K)^{\frac{1}{n-1}}\mathcal{E}_{0}^{-\frac{1}{n-1}}+\frac{1}{a}(2K)^{\frac{1}{n-1}}\mathcal{E}_{0}^{-\frac{1}{n-1}}\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}}).$
@@
$\mathcal{J}_{0}(\Theta)=\frac{1}{a}(2K)^{\frac{1}{n-1}}\Theta\mathcal{E}_{0}^{-\frac{1}{n-1}}+\frac{1}{a}(2K)^{\frac{1}{n-1}}\Theta\mathcal{E}_{0}^{-\frac{1}{n-1}}\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}}),$
where $\mathcal{E}_{0}=|v_{1}-v_{2}|^{2}$. ∎
In the above proof we used the fact that $\frac{\mathrm{d}u_{0}}{\mathrm{d}\mathcal{J}_{0}}=0$. In the following lemma we give the proof.
Lemma 3.5.
Proof.
With (3.9) we define
$F(u_{0},\mathcal{J}_{0}):=J_{0}^{2}\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{2}{n-1}}u_{0}^{2}+u_{0}^{n-1}-1.$
We observe that $F(1,0)=0$ and
$\frac{\partial F}{\partial u_{0}}(u_{0},\mathcal{J}_{0})=2J_{0}^{2}\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{2}{n-1}}u_{0}+(n-1)u_{0}^{n-2},$
so $\frac{\partial F}{\partial u_{0}}(1,0)=n-1\neq 0$, for $n\neq 1.$ Also
$\frac{\partial F}{\partial\mathcal{J}_{0}}(u_{0},\mathcal{J}_{0})=2J_{0}\bigg{(}K+\frac{\mathcal{E}_{0}}{2K}\bigg{)}^{\frac{2}{n-1}}u_{0}^{2},$
which is equivalent to
$\frac{\mathrm{d}u_{0}}{\mathrm{d}\mathcal{J}_{0}}=-\bigg{(}\frac{\partial F}{\partial u_{0}}(1,0)\bigg{)}^{-1}\frac{\partial F}{\partial\mathcal{J}_{0}}(1,0)=0,$
as required. ∎
We are ready to give the proof of the main proposition of this section which is about finding the formula (3.4) of the collision kernel $B$.
@@
$\mathcal{J}_{0}(\Theta)=\frac{(2K)^{\frac{1}{n-1}}}{a}\bigg{(}1+\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}})\bigg{)}\Theta\mathcal{E}_{0}^{-\frac{1}{n-1}},$
for any $n\in(2,5]$, for large relative velocities $|v_{1}-v_{2}|>0$ and for $\Theta\in(0,\frac{\pi}{2}-\sigma)$, $\sigma>0$, $a$ is a fixed real number and $K>0$, where $\mathcal{E}_{0}=|v_{1}-v_{2}|^{2}$. Then the derivative with respect to $\Theta$ is
$\partial_{\Theta}\mathcal{J}_{0}(\Theta)=\frac{(2K)^{\frac{1}{n-1}}}{a}\bigg{(}1+\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}})\bigg{)}\mathcal{E}_{0}^{-\frac{1}{n-1}}.$
@@
$\bigg{|}B$ $(v_{1}-v_{2},\Theta)-\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}|v_{1}-v_{2}|^{\frac{n-5}{n-1}}\bigg{|}$
$=\bigg{|}\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}\bigg{[}|v_{1}-v_{2}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})-|v_{1}-v_{2}|^{\frac{n-5}{n-1}}\bigg{]}\bigg{|}$
$=\bigg{|}\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})\bigg{|}$
$=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}}),$
@@
$\Theta=\Theta(\mathcal{E}_{0},\mathcal{J}_{0}):=\arcsin{\mathcal{J}_{0}}+\mathcal{J}_{0}\int_{\rho_{*}}^{1}\frac{\mathrm{d}\rho}{\rho^{2}\sqrt{1-\frac{4U(\rho)}{\mathcal{E}_{0}}-\frac{\mathcal{J}^{2}_{0}}{\rho^{2}}}},$
@@
$\Theta=\Theta(\mathcal{E}_{0},\mathcal{J}_{0})=\arcsin{\mathcal{J}_{0}}+\int_{\arcsin{\mathcal{J}_{0}}}^{\frac{\pi}{2}}\frac{\rho\mathcal{J}_{0}\mathcal{E}_{0}\sin\varphi}{\mathcal{E}_{0}\mathcal{J}_{0}^{2}-2\rho^{3}U^{\prime}(\rho)}\mathrm{d}\varphi$ (3.13)
@@
$\Theta=\Theta(\mathcal{E}_{0},\mathcal{J}_{0})=\arcsin{\mathcal{J}_{0}}+\int_{\arcsin{\mathcal{J}_{0}}}^{\frac{\pi}{2}}\frac{\rho\mathcal{J}_{0}\mathcal{E}_{0}\sqrt{\frac{4U(\rho)}{\mathcal{E}_{0}}+\frac{\mathcal{J}_{0}^{2}}{\rho^{2}}}}{\mathcal{E}_{0}\mathcal{J}_{0}^{2}-2\rho^{3}U^{\prime}(\rho)}\mathrm{d}\varphi.$
@@ 4.1. Energy estimates
[…]
$B(v_{1}-v_{2},\Theta)$ $=\frac{(2K)^{\frac{2}{n-1}}}{a^{2}}\left[|v_{1}-v_{2}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})\right]\Theta(\sin\Theta)^{-1},$
for $n\in(2,5]$, large $|v_{1}-v_{2}|>0$ and $\Theta\in(0,\frac{\pi}{2}-\sigma)$. Consider
@@
$=c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}|v|\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})B(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x$
$\quad-c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}|v|f_{t}(x,v)\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}g_{0}(\bar{v})B(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x,$
[…]
@@
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}D_{f}(t)$ $\leq c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}(|v^{\prime}|+|\bar{v}^{\prime}|)\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})B(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x$
$\quad-c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}|v|f_{t}(x,v)\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}g_{0}(\bar{v})B(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x$
@@
$\quad+c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|B(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x.$

### v2

@@ 3.2. Formula of the collision kernel $\mathcal{B}$
where the first term in the right hand side is a finite fixed number, since $n$ is fixed, so we define this finite fixed number by $a$. Furthermore, the second term in the right hand side is equal to $\mathcal{O}(\frac{\lambda}{b})$, as $\mathcal{E}_{0}\to\infty$. Therefore we have that
$\int_{0}^{1-\frac{\lambda}{b}}\frac{\mathrm{d}w}{\sqrt{1-(1-w)^{n-1}}}=a+\mathcal{O}(\frac{\lambda}{b}),\quad\text{as}\ \mathcal{E}_{0}\to\infty.$
Thus, we write
@@
$\Sigma^{\prime}(0)=1+a\bigg(1+\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{1}{n-1}}+\mathcal{O}(\frac{\lambda}{b})\bigg(1+\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{1}{n-1}}.$ (3.10)
Also, the second term of (3.9), for $u_{0}=1-\delta$, $\delta>0$ sufficient small, becomes
@@
$\int_{\bar{c}}^{1}\frac{\mathrm{d}y}{\sqrt{1-y^{n-1}}}=\int_{0}^{1-\bar{c}}\frac{\mathrm{d}z}{\sqrt{1-(1-z)^{n-1}}}=\int_{0}^{1-\bar{c}}\frac{\mathrm{d}z}{\sqrt{(n-1)z}}=\frac{2(1-\bar{c})^{\frac{1}{2}}}{\sqrt{n-1}}<\infty,$
where in the first equality we used the change of variable z=1yz=1-y and in the second equality we used Taylor expansion of first order of the function $(1-z)^{n-1}$ around $z_{0}=0$. Here, $\bar{c}$ is a positive fixed number, less than $1$. The function $j_{\delta}(y)$ can be written as
@@
$\int_{\frac{\lambda}{b(1-\delta)}}^{1}\frac{\mathrm{d}y}{\sqrt{j_{\delta}(y)}}\to\int_{\bar{c}}^{1}\frac{\mathrm{d}y}{\sqrt{1-y^{n-1}}}\quad\text{as}\ \delta\to 0.$
@@
$\lim_{\mathcal{E}_{0}\to\infty}\frac{\lambda}{b}\bigg(1+\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{1}{n-1}}=1,\quad\textrm{i.e.}\quad\mathcal{O}(\frac{\lambda}{b})\bigg(1+\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{1}{n-1}}=\mathcal{O}(1),\quad\text{as}\ \mathcal{E}_{0}\to\infty.$
Therefore
@@
$=a\bigg(\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{1}{n-1}}+\underbrace{\frac{a}{n-1}\bigg(\frac{\mathcal{E}_{0}}{2}\bigg)^{\frac{2-n}{n-1}}+\mathcal{O}(1)}_{=\mathcal{O}(1)},$
where in the second equality we used the Taylor expansion of first order with respect to $x$ around 0, evaluated at $x=2$ of the function $h(x):=\big(x+\mathcal{E}_{0}\big)^{\frac{1}{n-1}}$. Thus, by the inverse function theorem
$(\Sigma^{-1})^{\prime}(\Sigma(0))=\frac{1}{a\big(\frac{\mathcal{E}_{0}}{2}\big)^{\frac{1}{n-1}}+\mathcal{O}(1)}$ $=a^{-1}\bigg(\frac{\mathcal{E}_{0}}{2}\bigg)^{-\frac{1}{n-1}}\frac{1}{1+a^{-1}\big(\frac{\mathcal{E}_{0}}{2}\big)^{-\frac{1}{n-1}}\mathcal{O}(1)}$
$=a^{-1}\bigg(\frac{\mathcal{E}_{0}}{2}\bigg)^{-\frac{1}{n-1}}\big(1+\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}})\big),$
@@
$(\Sigma^{-1})^{\prime}(\Sigma(0))=\frac{2^{\frac{1}{n-1}}}{a}\mathcal{E}_{0}^{-\frac{1}{n-1}}+\frac{2^{\frac{1}{n-1}}}{a}\mathcal{E}_{0}^{-\frac{1}{n-1}}\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}}).$
@@
$\mathcal{J}_{0}(\Theta)=\frac{2^{\frac{1}{n-1}}}{a}\Theta\mathcal{E}_{0}^{-\frac{1}{n-1}}+\frac{2^{\frac{1}{n-1}}}{a}\Theta\mathcal{E}_{0}^{-\frac{1}{n-1}}\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}}),$
where $\mathcal{E}_{0}=|v_{1}-v_{2}|^{2}$. This completes the proof of the proposition. ∎
We are now ready to give the proof of the main proposition of this section which is about finding the formula (3.3) of the collision kernel $B$.
@@
$\mathcal{J}_{0}(\Theta)=\frac{2^{\frac{1}{n-1}}}{a}\bigg(1+\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}})\bigg)\Theta\mathcal{E}_{0}^{-\frac{1}{n-1}},$
for any $n\in(3,5]$, for large relative velocities $|v_{1}-v_{2}|>0$ and $a$ is a fixed real number, where $\mathcal{E}_{0}=|v_{1}-v_{2}|^{2}$. Then the derivative with respect to $\Theta$ is
$\partial_{\Theta}\mathcal{J}_{0}(\Theta)=\frac{2^{\frac{1}{n-1}}}{a}\bigg(1+\mathcal{O}(\mathcal{E}_{0}^{-\frac{1}{n-1}})\bigg)\mathcal{E}_{0}^{-\frac{1}{n-1}}.$
@@
$\bigg|\mathcal{B}$ $(v_{1}-v_{2},\Theta)-\frac{2^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}|v_{1}-v_{2}|^{\frac{n-5}{n-1}}\bigg|$
$=\bigg|\frac{2^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}\bigg[|v_{1}-v_{2}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})-|v_{1}-v_{2}|^{\frac{n-5}{n-1}}\bigg]\bigg|$
$=\bigg|\frac{2^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})\bigg|$
$=\frac{2^{\frac{2}{n-1}}}{a^{2}}\Theta(\sin\Theta)^{-1}\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}}),$
@@
[…]
@@
[…]
@@ 4.1. Energy estimates
To prove Theorem 1.1 we extend the work of [MST18] by looking at error estimates over longer times considering long-term dynamics. We are doing this by finding a quantitative error for the difference $|P_{t}(S)-\hat{P_{t}}(S)|$, detailed expressions are in Proposition 4.8 below. We first provide quantitative estimates for the momentum of the tagged particle, the expected number of collisions and estimates for recollisions. It will be enough to provide all these estimates for the idealised evolution. We will use that $\mathcal{B}$ is the collision kernel which by (3.3) has the form
$\mathcal{B}(v_{1}-v_{2},\Theta)$ $=\frac{2^{\frac{2}{n-1}}}{a^{2}}\left[|v_{1}-v_{2}|^{\frac{n-5}{n-1}}+\mathcal{O}(|v_{1}-v_{2}|^{\frac{n-7}{n-1}})\right]\Theta(\sin\Theta)^{-1},$
for $n\in(3,5]$ and large $|v_{1}-v_{2}|>0$ as well as its boundedness for bounded $|v_{1}-v_{2}|$ . Consider
@@
$=c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}|v|\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})\mathcal{B}(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x$
$\quad-c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}|v|f_{t}(x,v)\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}g_{0}(\bar{v})\mathcal{B}(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x,$
by integration by parts and by using the linear Boltzmann equation (1.3).
@@
By using the inequality (4.1), we get
$\frac{\mathrm{d}}{\mathrm{d}t}D_{f}(t)$ $\leq c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}(|v^{\prime}|+|\bar{v}^{\prime}|)\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})\mathcal{B}(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x$
$\quad-c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}|v|f_{t}(x,v)\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}g_{0}(\bar{v})\mathcal{B}(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x$
@@
$\quad+c\int_{\mathds{R}^{3}\times\mathds{R}^{3}}\int_{\mathbb{S}^{2}\times\mathds{R}^{3}}f_{t}(x,v^{\prime})g_{0}(\bar{v}^{\prime})|\bar{v}^{\prime}|\mathcal{B}(v-\bar{v},\omega)\,\mathrm{d}\bar{v}\,\mathrm{d}{\omega}\,\mathrm{d}v\,\mathrm{d}x.$